Given that Q= (3P-1)/2P+1,
(a) Find the value of Q when P=2, ( I solved and Q=1)
(b) Express P in terms of Q( Please help me with this!!!)

Answer by Theo(13342)

when p = 2, the formula becomes q = (3*2-1)/(2*2+1) which becomes 5/5 which is equal to 1.

expressing p in terms of q takes more work.

start with q = (3p-1)/(2p+1

multiply both sides of the eqution by (2p+1) to get:

q*(2p+1) = 3p-1

simplify by performing the indicated multiplication to get:

q*2p+q = 3p-1

subtract 3p from both ides of the equaqtion to get:

q*2p+q-3p = -1

subtract q from both sides of the equation to get:

q*2p-3p = -1-q

factor out the p on the left hand side of the equation to get:

p*(2q-3) = -1-q

divide both sides of the equation by (2q-3) to get:

p = (-1-q)/(2q-3)

-1-q can also be shown as -1(q+1) to get:

p = -(q+1)/(2q-3)

that's your equation.